OpenStudy (anonymous):

I'll medal someone if you explain and tell me the answer Find the product. 2/3 X (-1/2) A. -2/5 B. -1/3 C. 1/3 D. 2/5

OpenStudy (anonymous):

@LONGEST.NAME.POSSIBLE123456789

OpenStudy (anonymous):

@zelda101

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

negative x a positive is always a negative

OpenStudy (whpalmer4):

\[\frac{2}3*(-\frac{1}2)\] to multiply fractions, multiply the numerators to get the new numerator, and the denominators to get the new denominator

OpenStudy (anonymous):

So I eliminate C and D

OpenStudy (whpalmer4):

then count up the number of negative signs in front fractions. if that number is odd, put a negative sign in front of the answer.

OpenStudy (anonymous):

its b

OpenStudy (whpalmer4):

@zelda101 please do not give answers. openstudy is intended to guide students, not do the work for them.

OpenStudy (anonymous):

@whpalmer4 I gave you a medal

OpenStudy (anonymous):

im sorry

OpenStudy (whpalmer4):

@HalloimEllie so did you find the answer?

OpenStudy (anonymous):

@whpalmer4 Yes thank you for explaining it to me :)

OpenStudy (whpalmer4):

here's another one for you:\[(-\frac{1}2)(-\frac{1}3)=\]

OpenStudy (whpalmer4):

get it right and I will give you a medal 😀

OpenStudy (anonymous):

positive 1/6?

OpenStudy (whpalmer4):

my work here is done 😀

OpenStudy (anonymous):

Thank you!